A KMnO4 solution is standardized by a metallic iron primary standard. To do so, the metallic iron is dissolved in acid, reduced into Fe2+ and titrated by permanganate. If 33.00 mL of KMnO4 are required to titrate 0.5585 g of iron, what is the concentration of the permanganate solution?
Answer #1

The reaction is given as :

MnO4- + 8H+ + 5Fe2+ = Mn2+ + 4H2O + 5Fe3+

Number of moles of iron = 0.5585 / 55.85 = 0.01 mol

So number of moles of manganate required = 0.01 / 5 = 0.002 mol

Molarity = 0.002 / 0.033

= 0.06 M

So the concentartion of permanganate solution will be 0.06 M